OpenStudy (anonymous):

How do you find the value of y when 8^y = 2 ?

OpenStudy (lgbasallote):

change to log form \[\implies \log_8 2 = y\] now calculate in your calculator \(\log 8 \div \log 2\) to get your final answer does that help?

OpenStudy (anonymous):

or one can write\[(2^3)^y=2\]\[2^{3y}=2\]bases are equal then equate the exponents

OpenStudy (arlen):

@mukushla approach is quite fundamental ..go with it

OpenStudy (anonymous):

@Igbasallote Thank you :D